OpenStudy (anonymous):

How do I know where to start the graph when trying to graph something like g(x)=x2(squared) +3

OpenStudy (anonymous):

\(g = x^2 + 3\) ? What is the graph of \(g=x^2\)?

OpenStudy (anonymous):

where should the graph begin?

OpenStudy (anonymous):

What does the graph of \(x^2\) look like?

OpenStudy (anonymous):

I know that it's a parabola, but does it matter where it begins? Basically, does it have to be right at 0,0?

OpenStudy (anonymous):

am i correct?

OpenStudy (anonymous):

You are right parabola. if it says x^2 only it begins at 0,0. The + 3 is a hint. Move it up and start at (0,3).